Find the amount of silver liberated at cathode if 0.500 A of current is passed through AgNO base3 electrolyte for 1 hour. Atomic weight of silver is 107.9 g/mole.

Grade:upto college level

1 Answers

Deepak Patra
9 years ago
Sol. At Wt. At = 107.9 g/mole I = 0.500 A E base Ag = 107.9 g [As Ag is monoatomic] Z base Ag = E/f = 107.9/96500 = 0.001118 M = Zit = 0.001118 × 0.5 × 3600 = 2.01